OpenStudy (anonymous):

Find an equation of the circle whose diameter has endpoints (5,3) and (−3,1).

OpenStudy (anonymous):

The radius will be half the distance between these points. The center will be the midpoint.

OpenStudy (anonymous):

Once you get the radius \(r\) and the center \((h,k)\), then your equation is: \[ (x-h)^2+(y-k)^2=r^2 \]

OpenStudy (anonymous):

So would the center points = (1,2)? x= (5-3)/2 and y = (3+1)/2? Then use the radius formula \[\sqrt{(x2-x1)^{2}+(y 2-y1)^{2}}\] ? Or am I missing something

OpenStudy (zzr0ck3r):

correct